If Leah is 6 years older than Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue?



can you please show all steps. Thank you very much..

Answer by rfer(16322)

Set them all in terms of Sue's age as being x.
Sue=x
Leah=x+6
John=x+11
x+(x+6)+(x+11)=41
3x=24
x=8
Sue=8
Leah=14
John=19